OpenStudy (anonymous):

Is the line through points P(8, 3) and Q(4, 2) parallel to the line through points R(–5, –5) and S(–1, –2)? Yes; the lines have equal slopes. Yes; the lines are both vertical. No; the lines have unequal slopes. No; one line has zero slope and the other has no slope.

jimthompson5910 (jim_thompson5910):

Are you able to find the slopes of each line?

OpenStudy (anonymous):

i'm really bad at this so i dont know.

jimthompson5910 (jim_thompson5910):

Let's find the slope of the line through P and Q The point P is the first point (x1,y1), so (x1, y1) = (8, 3) This means x1 = 8, y1 = 3 The point Q is the first point (x2,y2), so (x2, y2) = (4, 2) This means x2 = 4, y2 = 2 Now use the slope formula m = (y2-y1)/(x2-x1) m = ( 2- 3)/(4-8) ... plug in the values of x1,y1,x2,y2 shown above m = (-1)/(-4) m = 1/4 So the slope of the line through P and Q is 1/4

jimthompson5910 (jim_thompson5910):

Now use the same basic steps to find the slope through R and S. I'll let you do this.

jimthompson5910 (jim_thompson5910):

oops typo i meant to write "The point Q is the SECOND point (x2,y2), "

jimthompson5910 (jim_thompson5910):

So are you able to find the slope of the line through R and S?